# A Malliavin calculus approach to score functions in diffusion generative models

### v3

@@ Abstract
Score-based diffusion generative models have recently emerged as a powerful tool for modelling complex data distributions. These models aim at learning the score function, which defines a map from a known probability distribution to the target data distribution via deterministic or stochastic differential equations (SDEs). The score function is typically estimated from data using a variety of approximation techniques, such as denoising or sliced score matching, Hyvärien’s method, or Schrödinger bridges. In this paper, we derive an exact, closed-form, expression for the score function for a broad class of nonlinear diffusion generative models. Our approach combines modern stochastic analysis tools such as Malliavin derivatives and their adjoint operators (Skorokhod integrals or Malliavin Divergence) with a new Bismut-type formula. The resulting expression for the score function can be written entirely in terms of the first and second variation processes, with all Malliavin derivatives systematically eliminated, thereby enhancing its practical applicability. The theoretical framework presented in this work offers a principled foundation for advancing score estimation methods in generative modelling, enabling the design of new sampling algorithms for complex probability distributions. Our results can be extended to broader classes of stochastic differential equations, opening new directions for the development of score-based diffusion generative models.
60H07,
60H10; 60J60; 68T07,
Score function,
Bismut formula,
Malliavin calculus,
stochastic differential equations,
generative modelling.,
@@
[class=MSC]
keywords:
[…]
and
@@ 1 Introduction
Historically, score-based generative models have been built upon diffusion processes described by linear Itô’s stochastic differential equations, i.e., SDEs with drift and diffusion coefficients that are linear in the state variables. Such a linearity assumption ensures analytical tractability and comes with other desirable properties. For example, the Fokker–Planck equation [7, 20] associated with the diffusion process admits a Gaussian stationary distribution and Gaussian transition densities, which have been used in the development of techniques such as Denoising Diffusion Probabilistic Models (DDPM) [9]. While the linearity assumption in diffusion processes has been instrumental in developing the first generation of score-based generative models, it also represents an overly restrictive simplification. Many real-world systems exhibit nonlinear dynamics, where both the drift and diffusion coefficients depend nonlinearly on the state variables. Extending diffusion models to such settings poses substantial challenges. Specifically, unlike linear SDEs, nonlinear SDEs typically lack closed-form expressions for their transition probability densities, making the computation of the score function $\nabla_{y}\log p_{t}(y)$ very challenging.
@@
The purpose of this paper is to bridge this gap by establishing a new rigorous link between score-based diffusion models and Malliavin calculus, a stochastic calculus of variations introduced by Paul Malliavin in the 1970s [11, 12]. Originally devised to investigate hypoelliptic operators and stochastic partial differential equations, Malliavin calculus provides a robust framework for analysing the smoothness of SDE solutions and their densities. By leveraging Malliavin derivatives and a novel Bismut-type formula we derive in this paper, which is distinct from the classical Bismut–Elworthy–Li formula [3, 5, 6] tailored to heat kernels, we derive an explicit (closed-form) representation of the score function for general nonlinear diffusion processes. The resulting formula is expressed entirely in terms of first and second variation processes [10, 2], with all Malliavin derivatives systematically eliminated. In particular, our framework generalises the state‐independent results of [14] to fully nonlinear SDEs with state‐dependent diffusion coefficients. For the numerical experiments of this framework, see the same work [14]. Very recently, other studies have applied Malliavin calculus to diffusion generative models for different ends: one focuses on robust numerical treatment of singular diffusion bridges and conditional control, and additionally derives an explicit yet distinct conditional-score formula via a generalised Tweedie identity [19], and another proposes an extension of diffusion models to infinite‐dimensional diffusion settings using Malliavin calculus [8].
@@
The implications of this new theoretical framework are profound. First, Malliavin calculus provides a powerful tool for overcoming the analytical challenges posed by nonlinear diffusion processes. Second, its rigorous foundations enable the development and analysis of new classes of generative models beyond linear diffusion. From a practical standpoint, this can potentially yield more accurate generative models for applications such as high-fidelity image synthesis, fluid dynamics [25], and protein folding simulations in molecular dynamics [1].
@@ 3 Malliavin calculus and covering vector fields
$\|u_{k}(t)\|^{2}$ $=\sum_{j,\ell}\left(\gamma^{-1}\right)_{jk}\left(\gamma^{-1}\right)_{\ell k}(D_{t}X_{T}^{j}\cdot D_{t}X_{T}^{\ell})\quad\text{(Euclidean inner product in $\mathbb{R}^{d}$)},$ (3.6)
@@
    $b_{n},b,\sigma_{n},\sigma\in C^{2}([0,T]\times\mathbb{R}^{m})$ with all first and second derivatives bounded uniformly in $n$, and $(b_{n},\sigma_{n})\to(b,\sigma)$ locally uniformly, and $\partial_{x}b_{n}\to\partial_{x}b,\partial_{x}\sigma_{n}\to\partial_{x}\sigma$ locally uniformly (i.e. $C^{1}_{\text{loc}}$-convergence).
@@
    For each $n$ and for $X$, the terminal value lies in $\mathbb{D}^{1,2}$, and their Malliavin covariance matrices
    $\gamma^{n}\;=\;\int_{0}^{T}D_{t}X^{n}_{T}\,(D_{t}X^{n}_{T})^{\top}\,dt,\quad\gamma\;=\;\int_{0}^{T}D_{t}X_{T}\,(D_{t}X_{T})^{\top}\,dt$
    satisfy $\gamma^{n},\gamma\geq\lambda I_{m}>0$ a.s. for some $\lambda>0$, together with $\sup_{n}\mathbb{E}[\|(\gamma^{n})^{-1}\|^{2}]<\infty$ and $\mathbb{E}[\|\gamma^{-1}\|^{2}]<\infty$.
@@
    The first variation processes $Y^{n}_{t}=\partial_{x}X^{n}_{t}$ satisfy $\sup_{n,t}\mathbb{E}[\|Y^{n}_{t}\|^{2}+\|(Y^{n}_{t})^{-1}\|^{2}]<\infty$, and $\sup_{n}\mathbb{E}\big{[}\sup_{t\in[0,T]}\|X^{n}_{t}\|^{2}\big{]}<\infty$ and $\mathbb{E}\big{[}\sup_{t\in[0,T]}\|X_{t}\|^{2}\big{]}<\infty$.
@@
    (Uniform $L^{4}$ Malliavin control) $\sup_{n}\mathbb{E}\!\int_{0}^{T}\|D_{t}X_{T}^{n}\|^{4}\,dt<\infty\quad\text{and}\quad\mathbb{E}\!\int_{0}^{T}\|D_{t}X_{T}\|^{4}\,dt<\infty.$
@@
Let $A_{n,R}=4T^{2}\sup_{(t,x)\in[0,T]\times B_{R}}|b_{n}(t,x)-b(t,x)|^{2}+4C_{\mathrm{BDG}}T\sup_{(t,x)\in[0,T]\times B_{R}}\|\sigma_{n}(t,x)-\sigma(t,x)\|^{2}$ and $K=4TC^{2}+4C_{\mathrm{BDG}}C^{2}$. By Gronwall’s inequality
@@ 4 Variation processes
For $t>T$, $D_{t}X_{T}=0$ (future perturbations don’t affect $X_{T}$). This follows because $D_{t}X_{s}=0$ for $s<t$, and $D_{t}X_{t}=\sigma(t,X_{t})$, with the perturbation propagating via $Y_{T,t}=Y_{T}Y_{t}^{-1}$.
@@ 5.1 Useful lemmas
We need to show $Y_{t}^{-1}$ is an Itô process. To this end, consider the SDE (5.1) with $Y_{0}=I_{m}$. The coefficients $\partial_{x}b(t,X_{t})$ and $\partial_{x}\sigma^{l}(t,X_{t})$ are assumed to be bounded and measurable. The SDE (5.1) is linear, and under standard conditions (e.g., Lipschitz continuity of the coefficients), the solution $Y_{t}$ exists, is unique, and is a semimartingale — specifically, an Itô process — adapted to the filtration generated by the Brownian motions $B_{t}^{l}$. Moreover, since $Y_{0}=I_{m}$ is invertible and the coefficients satisfy regularity conditions, $Y_{t}$ remains invertible almost surely for all $t\geq 0$. Define the function $f:\text{GL}(m,\mathbb{R})\to\text{GL}(m,\mathbb{R})$ by $f(A)=A^{-1}$, where $\text{GL}(m,\mathbb{R})$ is the group of $m\times m$ invertible matrices. The map $f$ is smooth (infinitely differentiable) on $\text{GL}(m,\mathbb{R})$, with first derivative $Df(A)H=-A^{-1}HA^{-1}$ and second derivative terms involving higher-order products. Applying Itô’ s formula to $Y_{t}^{-1}=f(Y_{t})$, where $Y_{t}$ is an Itô process, yields a stochastic differential of the form
@@
established in [15, Lemma 2.1.6], is justified. Without such a lower bound on the eigenvalues, the inverse may fail to be integrable. In dimension one, for instance, $A=\int_{0}^{T}B_{s}^{2}\,ds>0$ satisfies $A^{-q}\in L^{1}(\Omega)$ only for $q<\frac{1}{2}$; thus $A^{-1}\notin\mathbb{D}^{1,2}$ and Lemma 2.1.6 cannot be applied. Uniform ellipticity (or, more generally, the moment hypothesis $|\det A|^{-1}\in L^{p}$ for all $p\geq 1$ required in [15, Lemma 2.1.6]) prevents this pathology.
@@ 6 State-independent diffusion processes
In this section, we derive a simplified expression for the Skorokhod integral $\delta(u_{k})$ for SDEs with state-independent diffusion coefficients. Such SDEs are of interest in diffusion generative modelling because the noise amplitude remains identical across all sample paths, and can be chosen a priori via an appropriate noise scheduler. From a numerical standpoint, stochastic differential equations with state-independent diffusion coefficients exhibit reduced stiffness compared to those with multiplicative noise. This allows for larger time steps during sampling and offers direct control over the evolution of the marginal variance. This significantly simplifies the computation of likelihood weighting and score-matching losses. As a result, most practical architectures for diffusion-based generative modelling are based on state-independent diffusion coefficients – such as the VP (Variance-Preserving), VE (Variance-Exploding), and sub-VP SDEs introduced in [24]. Hence, consider SDEs of the form

### v4

@@ Abstract
Score-based diffusion generative models have recently emerged as a standard tool for modelling complex data distributions. These models aim at learning the score function, which defines a map from a known probability distribution to the target data distribution via deterministic or stochastic differential equations (SDEs). The score function is typically estimated from data using a variety of approximation techniques, such as denoising or sliced score matching, Hyvärien’s method, or Schrödinger bridges. In this paper, we derive an exact, closed-form expression for the score function for a broad class of nonlinear diffusion generative models. Our approach combines modern stochastic analysis tools such as Malliavin derivatives and their adjoint operators (Skorokhod integrals or Malliavin Divergence) with a new Bismut-type formula. The resulting expression for the score function can be written entirely in terms of the first and second variation processes, with all Malliavin derivatives systematically eliminated, thereby enhancing its practical applicability. The theoretical framework presented in this work offers a principled foundation for advancing score estimation methods in generative modelling, enabling the design of new sampling algorithms for complex probability distributions. Our results can be extended to broader classes of stochastic differential equations, opening new directions for the development of score-based diffusion generative models.
@@
Score function; Bismut formula; Malliavin calculus; stochastic differential equations; generative modelling
articletype: RESEARCH ARTICLE
@@ 1 Introduction
Historically, score-based generative models have been built upon diffusion processes described by linear Itô stochastic differential equations, i.e., SDEs with drift and diffusion coefficients that are linear in the state variables. Such a linearity assumption ensures analytical tractability and comes with other desirable properties. For example, the Fokker–Planck equation [7, 22] associated with the diffusion process admits a Gaussian stationary distribution and Gaussian transition densities, which have been used in the development of techniques such as Denoising Diffusion Probabilistic Models (DDPM) [9]. While the linearity assumption in diffusion processes has been instrumental in developing the first generation of score-based generative models, it also represents an overly restrictive simplification. Many real-world systems exhibit nonlinear dynamics, where both the drift and diffusion coefficients depend nonlinearly on the state variables. Extending diffusion models to such settings poses substantial challenges. Specifically, unlike linear SDEs, nonlinear SDEs typically lack closed-form expressions for their transition probability densities, making the computation of the score function $\nabla_{y}\log p_{t}(y)$ challenging.
@@
The purpose of this paper is to bridge this gap by establishing a new rigorous link between score-based diffusion models and Malliavin calculus, a stochastic calculus of variations introduced by Paul Malliavin in the 1970s [11, 12]. Originally devised to investigate hypoelliptic operators and Stochastic Partial Differential Equations (SPDEs), Malliavin calculus provides a rigorous framework for analysing the smoothness of SDE solutions and their densities. By leveraging Malliavin derivatives and a novel Bismut-type formula we derive in this paper, which is distinct from the classical Bismut–Elworthy–Li formula [3, 6, 5] tailored to heat kernels, we derive an explicit (closed-form) representation of the score function for general nonlinear diffusion processes. The resulting formula is expressed entirely in terms of first and second variation processes [10, 2], with all Malliavin derivatives systematically eliminated. In particular, our framework generalises the state‐independent results of [14] to fully nonlinear SDEs with state‐dependent diffusion coefficients. For the numerical experiments of this framework, see the same work [14]. Recently, complementary advances have appeared. First, an infinite-dimensional extension develops score-based diffusion modelling on separable Hilbert spaces using Malliavin tools and an operator-theoretic formulation [15]. A related direction employs Malliavin–Gamma calculus to treat random fields, providing an abstract Hilbertian setting for score-based diffusion generative models [8]. In parallel, a divergence–kernel representation is derived for scores of random dynamical systems, accommodating multiplicative noise and yielding additive-noise and short-time specialisations [16]. Finally, a path-space integration-by-parts approach addresses conditioning and diffusion bridges, producing a generalised Tweedie-type score identity and numerically robust training for singular rewards [21].
Beyond establishing the link with Malliavin calculus, our contribution is to move from gradient identities for expectations to score formulae that are usable in diffusion models. The central obstacle is that the classical Bismut–Elworthy–Li framework returns $\nabla_{x}\mathbb{E}[f(X_{T}^{x})]$ via Skorokhod (anticipative) integrals, whereas score-based methods require the density-level quantity $\nabla_{x}\log p_{t}(x)$. We address this by first replacing the anticipative objects with adapted, pathwise quantities, time integrals taken along the forward SDE, so that the resulting estimators are numerically implementable in both the linear regime (recovering the Fokker–Planck score) and the nonlinear setting with state-independent diffusion. Taken together, these steps convert the abstract BEL-type representations into practical, computationally robust algorithms for score-based diffusion modelling.
@@
The implications of this new theoretical framework are important. First, Malliavin calculus provides a tool for overcoming the analytical challenges posed by nonlinear diffusion processes. Second, its rigorous foundations enable the development and analysis of new classes of generative models beyond linear diffusion. From a practical standpoint, this can potentially yield more accurate generative models for applications such as high-fidelity image synthesis, fluid dynamics [27], and protein folding simulations in molecular dynamics [1].
@@ 3 Malliavin calculus and covering vector fields
$\|u_{k}(t)\|^{2}$ $=\sum_{j,\ell}\left(\gamma^{-1}\right)_{jk}\left(\gamma^{-1}\right)_{\ell k}(D_{t}X_{T}^{j}\cdot D_{t}X_{T}^{\ell})\quad\text{(Euclidean inner product)},$ (3.6)
@@
    We have $b_{n},b,\sigma_{n},\sigma\in C^{2}([0,T]\times\mathbb{R}^{m})$ with all first and second derivatives bounded uniformly in $n$, and $(b_{n},\sigma_{n})\to(b,\sigma)$ locally uniformly, and $\partial_{x}b_{n}\to\partial_{x}b$, $\partial_{x}\sigma_{n}\to\partial_{x}\sigma$ locally uniformly (i.e., $C^{1}_{\mathrm{loc}}$-convergence).
@@
    For each $n$ and for $X$, the terminal value lies in $\mathbb{D}^{1,2}$, and their Malliavin covariance matrices are given by $\gamma^{n}=\int_{0}^{T}D_{t}X_{T}^{n}(D_{t}X_{T}^{n})^{\top}\,dt$ and $\gamma=\int_{0}^{T}D_{t}X_{T}(D_{t}X_{T})^{\top}\,dt$, which satisfy $\gamma^{n},\gamma\geq\lambda I_{m}>0$ a.s. for some $\lambda>0$, together with $\sup_{n}\mathbb{E}[\|(\gamma^{n})^{-1}\|^{2}]<\infty$ and $\mathbb{E}[\|\gamma^{-1}\|^{2}]<\infty$.
@@
    The first variation processes $Y_{t}^{n}=\partial_{x}X_{t}^{n}$ satisfy the bounds $\sup_{n,t}\mathbb{E}[\|Y_{t}^{n}\|^{2}+\|(Y_{t}^{n})^{-1}\|^{2}]<\infty$, and we also have $\sup_{n}\mathbb{E}\big[\sup_{t\in[0,T]}\|X_{t}^{n}\|^{2}\big]<\infty$ and $\mathbb{E}\big[\sup_{t\in[0,T]}\|X_{t}\|^{2}\big]<\infty$.
@@
    (Uniform $L^{4}$ Malliavin control) We require that $\sup_{n}\mathbb{E}\int_{0}^{T}\|D_{t}X_{T}^{n}\|^{4}\,dt<\infty$ and $\mathbb{E}\int_{0}^{T}\|D_{t}X_{T}\|^{4}\,dt<\infty$.
@@
Let
$A_{n,R}$ $=4T^{2}\sup_{(t,x)\in[0,T]\times B_{R}}|b_{n}(t,x)-b(t,x)|^{2}$
$\quad+4C_{\mathrm{BDG}}T\sup_{(t,x)\in[0,T]\times B_{R}}\|\sigma_{n}(t,x)-\sigma(t,x)\|^{2}$ (3.9)
and $K=4TC^{2}+4C_{\mathrm{BDG}}C^{2}$. By Gronwall’s inequality
@@ 4 Variation processes
For $t>T$, $D_{t}X_{T}=0$ (future perturbations do not affect $X_{T}$). This follows because $D_{t}X_{s}=0$ for $s<t$, and $D_{t}X_{t}=\sigma(t,X_{t})$, with the perturbation propagating via $Y_{T,t}=Y_{T}Y_{t}^{-1}$.
@@ 5.1 Useful lemmas
We need to show $Y_{t}^{-1}$ is an Itô process. To this end, consider the SDE (5.1) with $Y_{0}=I_{m}$. The coefficients $\partial_{x}b(t,X_{t})$ and $\partial_{x}\sigma^{l}(t,X_{t})$ are assumed to be bounded and measurable. The SDE (5.1) is linear, and under standard conditions (e.g., Lipschitz continuity of the coefficients), the solution $Y_{t}$ exists, is unique, and is a semimartingale, specifically, an Itô process, adapted to the filtration generated by the Brownian motions $B_{t}^{l}$. Moreover, since $Y_{0}=I_{m}$ is invertible and the coefficients satisfy regularity conditions, $Y_{t}$ remains invertible almost surely for all $t\geq 0$. Define the function $f:\text{GL}(m,\mathbb{R})\to\text{GL}(m,\mathbb{R})$ by $f(A)=A^{-1}$, where $\text{GL}(m,\mathbb{R})$ is the group of $m\times m$ invertible matrices. The map $f$ is smooth (infinitely differentiable) on $\text{GL}(m,\mathbb{R})$, with first derivative $Df(A)H=-A^{-1}HA^{-1}$ and second derivative terms involving higher-order products. Applying Itô’ s formula to $Y_{t}^{-1}=f(Y_{t})$, where $Y_{t}$ is an Itô process, yields a stochastic differential of the form
@@
established in [17, Lemma 2.1.6], is justified. Without such a lower bound on the eigenvalues, the inverse may fail to be integrable. In dimension one, for instance, $A=\int_{0}^{T}B_{s}^{2}\,ds>0$ satisfies $A^{-q}\in L^{1}(\Omega)$ only for $q<\frac{1}{2}$; thus $A^{-1}\notin\mathbb{D}^{1,2}$ and Lemma 2.1.6 cannot be applied. Uniform ellipticity (or, more generally, the moment hypothesis $|\det A|^{-1}\in L^{p}$ for all $p\geq 1$ required in [17, Lemma 2.1.6]) prevents this pathology.
@@ 6 State-independent diffusion processes
In this section, we derive a simplified expression for the Skorokhod integral $\delta(u_{k})$ for SDEs with state-independent diffusion coefficients. Such SDEs are of interest in diffusion generative modelling because the noise amplitude remains identical across all sample paths, and can be chosen a priori via an appropriate noise scheduler. From a numerical standpoint, stochastic differential equations with state-independent diffusion coefficients exhibit reduced stiffness compared to those with multiplicative noise. This allows for larger time steps during sampling and offers direct control over the evolution of the marginal variance. This simplifies the computation of likelihood weighting and score-matching losses. As a result, most practical architectures for diffusion-based generative modelling are based on state-independent diffusion coefficients, such as the VP (Variance-Preserving), VE (Variance-Exploding), and sub-VP SDEs introduced in [26]. Hence, consider SDEs of the form
@@
Acknowledgements
AI assistance. We used ChatGPT only for minor editing (typos and wording). No proofs or mathematical content were generated, and no scientific content was altered.
Disclosure statement
The authors declare no potential conflict of interest.